# One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

### v2

@@ Abstract
Language is not monolithic. While benchmarks, including those designed for multiple languages, are often used as proxies to evaluate the performance of Large Language Models (LLMs), they tend to overlook the nuances of within-language variation, and thus fail to model the experience of speakers of non-standard dialects. Focusing on African American Vernacular English (AAVE), we present the first study aimed at objectively assessing the fairness and robustness of LLMs in handling dialects in canonical reasoning tasks, including algorithm, math, logic, and integrated reasoning. We introduce ReDial (Reasoning with Dialect Queries), a benchmark containing 1.2K+ parallel query pairs in Standardized English and AAVE. We hire AAVE speakers, including experts with computer science backgrounds, to rewrite seven popular benchmarks, such as HumanEval and GSM8K. With ReDial, we evaluate widely used LLMs, including GPT, Claude, Llama, Mistral, and the Phi model families. Our findings reveal that almost all of these widely used models show significant brittleness and unfairness to queries in AAVE. Our work establishes a systematic and objective framework for analyzing LLM bias in dialectal queries. Moreover, it highlights how mainstream LLMs provide unfair service to dialect speakers in reasoning tasks, laying a critical foundation for relevant future research. [^1] Code and data can be accessed here.
One Language, Many Gaps: Evaluating Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks
Fangru Lin1thanks: Work done during internship at Microsoft Research. Contact: fangru.lin@ling-phil.ox.ac.uk, shaoguang.mao@microsoft.com, Shaoguang Mao2, Emanuele La Malfa1,3, Valentin Hofmann4,5 Adrian de Wynter6,7, Xun Wang6, Si-Qing Chen6 Michael Wooldridge1,3, Janet B. Pierrehumbert1, Furu Wei2 1University of Oxford, 2Microsoft Research, 3Alan Turing Institute, 4Allen Institute for AI, 5University of Washington, 6Microsoft Corporation, 7University of York
@@ 1 Introduction
Figure 1: ReDial is a dialect reasoning benchmark composed of $1,200$+ SE-AAVE parallel queries. Its source data comes from existing benchmarks in SE. AAVE speakers are hired to rewrite each instance in their dialect but preserve their original intent, meaning, and ground truth output label to form their AAVE counterparts.
However, existing dialectal benchmarks (e.g., Ziems et al., 2023) do not cover these tasks, and current popular reasoning benchmarks such as HumanEval (Chen et al., 2021) and GSM8K (Cobbe et al., 2021) are constructed in Standardized English (SE). This could disadvantage dialect speakers in real-world applications like educational assessment (González-Calatayud et al., 2021), personalized recommendation (Kantharuban et al., 2024), and even multimodal tasks (e.g., voice assistant) Martin and Wright (2022), ultimately forcing them to shift their language styles to SE (Cunningham et al., 2024) in order to access the full benefits of modern technologies.
We present the first study that systematically and objectively evaluates LLM fairness and robustness in reasoning tasks expressed in a non-standard dialect. We choose AAVE since around $33$ million people worldwide and approximately $80\%$ of African Americans in the United States speak AAVE, with reports of discriminative behaviors in various scenarios (Lippi-Green, 1997; Purnell et al., 1999; Massey and Lundy, 2001; Grogger, 2011; Rickford and King, 2016). We hire AAVE speakers to manually rewrite instances from seven popular SE reasoning benchmarks into AAVE (Section 2.1). Our approach has unique advantages compared with prior works which either (i) rely on predefined lexical or morphosyntactic transformation rules (Ziems et al., 2022, 2023), which may overlook subtle contextual nuances, or (ii) use LLM as translators (Gupta et al., 2024), which may have the very biases that our research wants to unveil (Fleisig et al., 2024; Smith et al., 2024).
Category Algorithm (25.7%) Logic (29.8%) Math (24.7%) Intergrated (19.7%) Total
@@
Table 1: ReDial contains $1,216$ fully-annotated parallel prompts for four categories, drawn from seven data sources. Each category contribution to the total amount is reported in percentage points in brackets.
@@
Using ReDial, we benchmark widely used LLMs, including GPT-o1, GPT-4o, Claude-3.5-Sonnet, LLaMA-3.1-70B-Instruct, and others (Section 3). We find that almost all models experience statistically significant performance drops on AAVE prompts, despite their semantic equivalence to their SE counterparts. On average, we observe a relative performance reduction of more than 10%. This discrepancy persists even with advanced prompting techniques like Chain-of-Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022), indicating that current LLMs are both brittle and unfair with dialectal inputs.
@@ 2 Dataset
Following Zhu et al. (2023a), ReDial includes four canonical reasoning tasks—algorithm, logic, math. We additionally consider integrated reasoning as a compositional task requiring multiple skills. The task formulation of ReDial is linguistically diverse, addresses cornerstone problems in human reasoning, and is of particular interest as it is challenging for LLMs.
@@ 2.1 Data Sourcing
We construct a highly curated dataset by drawing upon seven established benchmarks covering different aspects of reasoning. For each source, we provide key references, task descriptions, and sample sizes. Additional examples can be found in Appendix A.1.
Figure 2: Annotation and cross-validation of ReDial instances. We first sample instances from datasets of four canonical reasoning tasks to compose the source data, then we hire AAVE speakers to rewrite the instances in their dialect. To ensure the quality, we conduct naturalness check by AAVE speakers and correctness check by non-AAVE speakers and LLMs. We reannotate instances that do not pass the quality checks and iterate the process until the data meet our criteria. Finally, we combine the source data and AAVE rewriting to obtain ReDial.
@@
Intergrated AsyncHow (Lin et al., 2024) is a planning reasoning benchmark. LLMs must interpret natural language descriptions (i.e., logic), find different possible paths in the graph (i.e., algorithm), and then calculate and compare the time cost for these paths (i.e., math) to reach the correct answer. We use stratified sampling based on the dataset’s complexity metrics and include $240$ instances.
@@ 2.2 Annotation and Quality Assurance
After data sourcing, we hire AAVE speakers to rewrite each instance in AAVE. We schematize our annotation and validation in Figure 2 and describe them below.
@@
We hire and ask AAVE speakers to rewrite each SE query so that it sounds natural to AAVE speakers while retaining all critical information (e.g., numerical values, logical conditions, and technical details). For algorithm-related tasks involving code, we hire annotators with a background in computer science to keep the logic and semantics of the code tasks. [^3] Details on annotator compensation, qualifications, and guidelines are presented in Appendix A.3.
@@
We then perform a careful quality check to ensure both naturalness and correctness. First, we ask annotators to cross-check and edit each others’ annotations to make sure that the annotations are natural to AAVE speakers. Second, to ensure correctness, we first have non-AAVE speakers manually check the essential information, then conduct a sanity check with GPT-4o for the correctness of rewriting (details in Appendix A.4). We manually check data that GPT-4o flags as invalid to see if all essential information is preserved. No instance is rejected solely based on the LLM’s judgment. We return invalid instances to AAVE speakers for correction and iterate the process until all data pass the check.
@@ 3.1 Experimental Setting
3.1.1 Models
GPT.
We use GPT-o1 (OpenAI, ), GPT-4o (OpenAI, 2024), GPT-4 (Achiam et al., 2023), GPT-3.5-turbo (Achiam et al., 2023), [^4] Proprietary model API version information: o1: GPT-o1-preview; gpt-4o: 2024-05-13; gpt-4: 2024-05-03; gpt-3.5-turbo: 2023-11-06. as a family of closed-source models to compare with open-source models for dialect robustness. In particular, o1 is trained using large-scale reinforcement learning (RL) to reason through a chain of thoughts and scales inference time computation to achieve highly complex reasoning paths. It has demonstrated significant improvements in reasoning tasks (OpenAI, ). We use GPT-o1 model to understand how RL reasoning post-training affects LLMs’ dialect robustness and fairness.
Claude.
Developed by Anthropic, the Claude 3 model family represents a widely-used proprietary LLM. For our experiments, we utilize the Claude 3.5 Sonnet model  (Anthropic, ).
LLaMA.
Mistral / Mixtral.
We use Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Mistral-7B-Instruct-v0.3 is reported to be outstanding in reasoning; with Mixtral-8x7B-Instruct-v0.1, we can understand whether Mixture-of-Expert (MoE) architectures enhance dialect robustness.
Phi.
We use Phi-3-Mini / Small / Medium-128K-Instruct (Abdin et al., 2024; Gunasekar et al., 2023) in our experiment. Phi-3 models, pre-trained on carefully designed “textbook” data, are reported for impressive performance in reasoning despite their small sizes (3.8/7/14B parameters each). We use these models to understand how highly curated pre-training data affect LLMs’ dialect robustness and fairness.
@@ 3.1.2 Implementation and Evaluation
We set the temperature to zero for the main experiments to ensure maximum reproducibility. We report two prompting methods in our main results: (i) direct prompting LLMs with task instances, which resembles general real-life use cases the most (Direct) and (ii) zero-shot Chain of Thought (Wei et al., 2022; Kojima et al., 2022) (CoT, i.e., adding instructions in the spirit of “Let’s think step by step” on top of task descriptions, which resembles expert user prompts to improve model performance). For GPT-o1, we only test the direct prompting as its inherent CoT reasoning pattern. [^5] We also test non-zero temperatures and report results in Appendix A.6. We report further implementation details in Appendix A.5.666We deliberately avoid testing advanced prompting methods, such as Tree of Thought (Yao et al., 2024) and Self-Refine (Madaan et al., 2024). Our focus is on evaluating how LLMs perform when prompted for everyday use by dialect users, which is critical for assessing fairness in LLMs. Similarly, we do not fine-tune any models, as our study aims to investigate biases inherent in base models. The effects of fine-tuning are beyond the scope of this study.
Algorithm Math Logic Intergrated All
@@
Table 3: Pass rates by task averaged across responses from all models with direct prompting. In bold, results statistically deviate according to McNemar’s tests applied to AAVE and SE (i.e., models have significant drops in AAVE). We also report the SAE-AAVE absolute delta in performance.
@@ 3.2 Experimental Results
The absolute performance gaps commonly range from around 5% to over 10% ($\Delta$ in table 2). Specifically, GPT-4o (zero-shot) shows an absolute gap of 11.6%, dropping from an average of 0.832 to 0.716. GPT-4 (CoT) exhibits an 11.5% drop. Mixtral-8x7B-Instruct-v0.1 (zero-shot) shows a particularly large difference of 11.4% points as well. Interestingly, we found that although the performance drop of GPT-o1 is smaller than other GPT models, but still significant. It indicates that although further RL post-training on general reasoning and inference scaling can systematically enhance dialect robustness and fairness, they cannot completely solve the problem.
In short, dialect unfairness and brittleness are identified in all the models we examined, including the MoE and large reasoning models. This finding indicates that the problem is widespread, non-trivial, and cannot be easily mitigated by naively changing model architecture or proposing more complex reasoning paths.
@@
When aggregated by task type, AAVE queries cause a statistically significant performance drop across all these categories (Table 3). For instance, when averaging results across all models: direct prompting leads to an average 10% relative performance drop.
@@
While CoT prompting can slightly reduce the discrepancy for some models, it fails to close it entirely. For example, GPT-4o’s performance gap decreases from about 0.116 (zero-shot) to 0.055 (CoT). This suggests that even when models are given additional reasoning "scaffolding," their understanding and performance in AAVE remain comparatively weaker than in SE, which is also in line with our observation with GPT-o1 results.
@@ 4 Discussions
This section investigates the potential reasons for AAVE’s brittleness. We show that LLMs’ brittleness with AAVE reasoning queries cannot be simply attributed to the lack of understanding of this dialect or simple lexical features. The nuanced conversational norms of AAVE also play an important role in the problems for LLMs.
@@ 4.1 General Understanding and Morphosyntactic Features
One possible explanation for the performance drop is that LLMs cannot process AAVE dialect. We thus computed perplexities on ReDial AAVE vs. SAE prompts. Indeed, Table 4 confirms that LLMs exhibit higher perplexities on dialect than SE.
@@
However, is the insufficient understanding the only reason that is causing LLMs’ performance to drop? To answer it, we gradually inject typos in SE ReDial by replacing/deleting/adding words/characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We find that while these perturbations degrade model performance, the drop does not reach the severity observed with natural AAVE data on large-scale models (see full results in Section A.8). This discrepancy suggests that AAVE brittleness is not solely due to the general difficulty in processing. There are some dialect-specific reasons for the performance drop.
If language-agnostic processing ability cannot explain LLMs’ brittleness, can we attribute the problem to morphosyntactic AAVE features? Following Ziems et al. (2022, 2023), we use morphosyntactic transformation rules to inject AAVE features into SE ReDial. We find that performance degradation generally intensifies as the density of AAVE-specific features increases when we gradually inject AAVE features into standard ReDial (see full results in Section A.9). This suggests that these features play a significant role in diminishing model performance.

### v3

# Assessing Dialect Fairness and Robustness of Large Language Models in Reasoning Tasks

@@ Abstract
Language is not monolithic. While benchmarks, including those designed for multiple languages, are often used as proxies to evaluate the performance of Large Language Models (LLMs), they tend to overlook the nuances of within-language variation and thus fail to model the experience of speakers of non-standard dialects. Focusing on African American Vernacular English (AAVE), we present the first study aimed at objectively assessing the fairness and robustness of LLMs in handling dialects across canonical reasoning tasks, including algorithm, math, logic, and integrated reasoning. We introduce ReDial (Reasoning with Dialect Queries), a benchmark containing 1.2K+ parallel query pairs in Standardized English and AAVE. We hire AAVE speakers, including experts with computer science backgrounds, to rewrite seven popular benchmarks, such as HumanEval and GSM8K. With ReDial, we evaluate widely used LLMs, including GPT, Claude, Llama, Mistral, and the Phi model families. Our findings reveal that almost all of these widely used models show significant brittleness and unfairness to queries in AAVE. Our work establishes a systematic and objective framework for analyzing LLM bias in dialectal queries. Moreover, it highlights how mainstream LLMs provide unfair service to dialect speakers in reasoning tasks, laying a critical foundation for future research. [^1] Code and data can be accessed here.
Assessing Dialect Fairness and Robustness of
Large Language Models in Reasoning Tasks
Fangru Lin1thanks: This work was conducted during an internship at Microsoft Research. Contact: fangru.lin@ling-phil.ox.ac.uk, dawnmsg941107@gmail.com  Shaoguang Mao2  Emanuele La Malfa1,3  Valentin Hofmann4,5 Adrian de Wynter6,7  Xun Wang6  Si-Qing Chen6 Michael Wooldridge1,3  Janet B. Pierrehumbert1  Furu Wei2 1University of Oxford  2Microsoft Research  3The Alan Turing Institute 4Allen Institute for AI  5University of Washington 6Microsoft Corporation  7University of York
@@ 1 Introduction
Figure 1: ReDial is a dialect reasoning benchmark composed of 1.2K+ SE-AAVE parallel queries. Its source data comes from existing benchmarks in SE. AAVE speakers are hired to rewrite each instance in their dialect but preserve their original intent, meaning, and ground truth output label to form their AAVE counterparts. Highlighted parts in blue are major differences in AAVE rewriting compared to SE.
However, existing dialectal benchmarks (e.g., Ziems et al., 2023) do not cover these tasks, and current popular reasoning benchmarks such as HumanEval (Chen et al., 2021) and GSM8K (Cobbe et al., 2021) are constructed in Standardized English (SE). This could disadvantage dialect speakers in real-world applications like educational assessment (González-Calatayud et al., 2021), personalized recommendation (Kantharuban et al., 2024), and even multimodal tasks (e.g., voice assistants; Martin and Wright, 2022), ultimately forcing them to shift their language styles to SE (Cunningham et al., 2024) in order to access the full benefits of modern technologies, even though they prefer to use their own dialects Blaschke et al. (2024).
We present the first study that systematically and objectively evaluates LLM fairness and robustness in reasoning tasks expressed in a non-standard dialect. We choose AAVE since around $33$ million people worldwide and approximately $80\%$ of African Americans in the United States speak AAVE, with reports of discriminative behaviors in various scenarios (Lippi-Green, 1997; Purnell et al., 1999; Massey and Lundy, 2001; Grogger, 2011). Moreover, relevant research shows that many speakers of English have difficulty understanding AAVE Rickford and King (2016). With the objective of making sure that users can use the language style they prefer instead of being constrained by the preference of the language model service, we consider that there is a need to separately consider AAVE from SE in evaluating LLMs. We hire AAVE speakers to manually rewrite instances from seven popular SE reasoning benchmarks into AAVE (Section 2.1). Our approach has unique advantages compared with prior works that either (i) rely on predefined lexical or morphosyntactic transformation rules (Ziems et al., 2022, 2023), which may overlook subtle contextual nuances, or (ii) use LLMs as translators (Gupta et al., 2024), which may have the very biases that our research wants to unveil (Fleisig et al., 2024; Smith et al., 2024).
Category Algorithm (25.7%) Logic (29.8%) Math (24.7%) Integrated (19.7%) Total
@@
Table 1: ReDial contains 1,216 fully-annotated parallel prompts for four categories, drawn from seven data sources. Each category’s contribution to the total is reported in percentage points in brackets.
@@
Using ReDial, we benchmark widely used LLMs, including GPT-o1, GPT-4o, Claude-3.5-Sonnet, Llama-3.1-70B-Instruct, and others (Section 3). We find that almost all models experience statistically significant performance drops on AAVE prompts, despite their semantic equivalence to their SE counterparts. On average, we observe a relative performance reduction of more than 10%. This discrepancy persists even with advanced prompting techniques like Chain of Thought (CoT) prompting (Kojima et al., 2022; Wei et al., 2022), indicating that current LLMs are both brittle and unfair with dialectal inputs.
@@ 2 Dataset
Following Zhu et al. (2023a), ReDial includes canonical reasoning tasks—algorithm, logic, and math. We additionally consider integrated reasoning as a compositional task requiring multiple skills. The task formulation of ReDial is linguistically diverse, addresses cornerstone problems in human reasoning, and is of particular interest as it is challenging for LLMs.
@@ 2.1 Data Sourcing
We construct a highly curated dataset by drawing upon seven established benchmarks covering different aspects of reasoning. We purposefully select the benchmarks that can capture the needs of real-world applications.
For each source, we provide key references, task descriptions, and sample sizes. Additional examples can be found in Appendix A.1.
Figure 2: Annotation and cross-validation of ReDial instances. We first sample instances from datasets of four canonical reasoning tasks to compose the source data, then we hire AAVE speakers to rewrite the instances in their dialect. To ensure the quality, we conduct a naturalness check by AAVE speakers and a correctness check by non-AAVE speakers and LLM. We reannotate instances that do not pass the quality checks and iterate the process until the data meets our criteria. Finally, we combine the source data and AAVE rewriting to obtain ReDial.
@@
Integrated AsyncHow (Lin et al., 2024) is a planning reasoning benchmark. LLMs must interpret natural language descriptions (i.e., logic), find different possible paths in the graph (i.e., algorithm), and then calculate and compare the time cost for these paths (i.e., math) to reach the correct answer. We use stratified sampling based on the dataset’s complexity metrics and include 240 instances.
@@ 2.2 Annotation and Quality Assurance
After data sourcing, we hire AAVE speakers to rewrite each instance in AAVE. We schematize our annotation and validation in Figure 2 and describe them below, by which we ensure the consistency, representativeness, and neutrality of our dataset.
@@
We hire and instruct AAVE speakers to rewrite each SE query so that it sounds natural to AAVE speakers while retaining all critical information (e.g., numerical values, logical conditions, and technical details). For algorithm-related tasks involving code, we hire annotators with a background in computer science to ensure the logic and semantics of the code tasks. In total, we hire 13 annotators with different demographic backgrounds to reduce personal biases. [^3] Details on annotator compensation, qualifications, guidelines, and demographic distribution are presented in Appendix A.3.
@@
We then perform a careful quality check to ensure both naturalness and correctness. First, we ask annotators to cross-check and edit each others’ annotations to make sure that the annotations are natural to AAVE speakers. This can also further reduce individual annotator bias. Second, to ensure correctness, we first have non-AAVE speakers manually check the essential information, then conduct a sanity check with GPT-4o for the correctness of rewriting (details in Appendix A.4). We manually check data that GPT-4o flags as invalid to see if all essential information is preserved. No instance is rejected solely based on the LLM’s judgment. We return invalid instances to AAVE speakers for correction and iterate the process until all data passes the check.
@@ 3.1 Experimental Setting
3.1.1 Models
GPT.
We use GPT-o1 (OpenAI, 2024b), GPT-4o (OpenAI, 2024a), GPT-4 (Achiam et al., 2023), GPT-3.5-turbo (Achiam et al., 2023), [^4] Proprietary model API version information: o1: GPT-o1-preview; gpt-4o: 2024-05-13; gpt-4: 2024-05-03; gpt-3.5-turbo: 2023-11-06. as a family of closed-source models to compare with open-source models for dialect robustness. In particular, o1 is trained using large-scale reinforcement learning (RL) to reason through CoT and scales inference time computation to achieve highly complex reasoning paths, demonstrating significant improvements in reasoning tasks (OpenAI, 2024b). We use the GPT-o1 model to understand how RL reasoning post-training affects LLMs’ dialect robustness and fairness.
Claude.
Developed by Anthropic, the Claude 3 model family represents a widely-used proprietary LLM. For our experiments, we utilize the Claude 3.5 Sonnet model (Anthropic, 2024).
Llama.
Mistral / Mixtral.
We use Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024). Mistral-7B-Instruct-v0.3 is reported to be outstanding in reasoning; with Mixtral-8x7B-Instruct-v0.1, we can understand whether Mixture-of-Expert architectures enhance dialect robustness.
Phi.
We use Phi-3-Mini / Small / Medium-128K-Instruct (Gunasekar et al., 2023; Abdin et al., 2024) in our experiment. Phi-3 models, pre-trained on carefully designed “textbook” data, are reported for impressive performance in reasoning despite their small sizes (3.8/7/14B parameters each). We use these models to understand how highly curated pre-training data affect LLMs’ dialect robustness and fairness.
Figure 3: Model pass rate and average response token count before and after being prompted for standardization. Standardization prompting generally improves LLM performance in both SE and AAVE ReDial (bar plot). However, even AAVE ReDial with standardization prompting cannot reach LLMs’ vanilla performance in SE ReDial, even though they also tend to result in more tokens generated and thus higher inference cost (scatter plot).
@@ 3.1.2 Implementation and Evaluation
We set the temperature to zero for the main experiments to ensure maximum reproducibility. We report two prompting methods in our main results: (i) direct prompting LLMs with task instances, which resembles general real-life use cases the most (Direct) and (ii) zero-shot Chain of Thought (CoT; Kojima et al., 2022; Wei et al., 2022), i.e., adding instructions in the spirit of “Let’s think step by step” on top of task descriptions, which resembles expert user prompts to improve model performance. For GPT-o1, we only test direct prompting due to its inherent CoT reasoning pattern. [^5] We also test non-zero temperatures for a subset of the models and report results in Appendix A.6. We report further implementation details in Appendix A.5.666We deliberately avoid testing advanced prompting methods, such as Tree of Thought (Yao et al., 2024) and Self-Refine (Madaan et al., 2024). Our focus is on evaluating how LLMs perform when prompted for everyday use by dialect users, which is critical for assessing fairness in LLMs. Similarly, we do not fine-tune any models, as our study aims to investigate biases inherent in models prior to task-specific adaptation. The effects of fine-tuning are beyond the scope of this study.
Algorithm Math Logic Integrated All
@@
Table 3: Pass rates by task averaged across responses from all models with direct prompting. In bold, results show statistically significant differences according to McNemar’s tests applied to AAVE and SE (i.e., models have significant drops in AAVE). We also report the SE-AAVE absolute delta in performance.
@@ 3.2 Experimental Results
The absolute performance gaps commonly range from around 5% to over 10% ($\Delta$ in Table 2). Specifically, GPT-4o (zero-shot) shows an absolute gap of 11.6%, dropping from an average of 0.832 to 0.716. GPT-4 (CoT) exhibits an 11.5% drop. Mixtral-8x7B-Instruct-v0.1 (zero-shot) shows a particularly large difference of 11.4% points as well. Interestingly, we found that although the performance drop of GPT-o1 is smaller than other GPT models, it is still significant. This indicates that although further RL post-training on general reasoning and inference scaling can systematically enhance dialect robustness and fairness, they cannot completely solve the problem.
In short, dialect unfairness and brittleness are identified in all the models we examined, including the mixture of expert and large reasoning models. This finding indicates that the problem is widespread, non-trivial, and cannot be easily mitigated by naively changing model architecture or proposing more complex reasoning paths.
@@
When aggregated by task type, AAVE queries cause a statistically significant performance drop across all these categories (Table 3). For instance, when averaging results across all models, inputs written in AAVE (via direct prompting) lead to an average 10% relative performance drop.
@@
While CoT prompting can slightly reduce the discrepancy for some models, it fails to close it entirely. For example, GPT-4o’s performance gap decreases from about 0.116 (zero-shot) to 0.043 (CoT). This suggests that even when models are given additional reasoning “scaffolding,” their understanding and performance in AAVE remain comparatively weaker than in SE, which is also in line with our observation with GPT-o1 results. We also try to bridge the gap by telling the LLMs to rephrase in Standard English then answer the question (i.e., standardization), but this does not cancel the performance gap, while only introducing more inference cost (Figure 3). This means that even if dialect users pay more, they might still not receive the same quality service as users who use SE.
@@ 4 Discussions
This section investigates the potential reasons for AAVE’s brittleness. We show that LLMs’ brittleness with AAVE reasoning queries is not simply due to the lack of understanding of this dialect or simple lexical features. The nuanced conversational norms of AAVE also contribute significantly to LLMs’ difficulties.
@@ 4.1 General Understanding and Morphosyntactic Features
One possible explanation for the performance drop is that LLMs cannot process AAVE. We thus computed perplexities on ReDial AAVE vs. SE prompts in Llama-3.1-70B-Instruct, Phi-3-Medium-128K-Instruct, and Phi-3-Mini-128K-Instruct. Indeed, Table 4 confirms that LLMs exhibit higher perplexities on dialect than SE.
@@
However, is the insufficient understanding the only reason for LLMs’ performance to drop? To investigate this further, we gradually inject typos into SE ReDial by replacing/deleting/adding words and characters, such that we make the input texts more difficult for LLMs (i.e., the measured perplexity goes up). We present results in Figure 4. While these perturbations degrade model performance, the drop does not reach the severity observed with natural AAVE data on large-scale models. This suggests that AAVE brittleness is not solely due to difficulties in text comprehension.
If language-agnostic processing ability cannot explain LLMs’ brittleness, can we attribute the problem to morphosyntactic AAVE features? Following Ziems et al. (2022, 2023), we use morphosyntactic transformation rules to inject AAVE features into SE ReDial. We find that performance degradation generally intensifies as the density of AAVE-specific features increases (see full results in Appendix A.7). This suggests that these features play a significant role in diminishing model performance.